Suppose that twenty-eight percent of women consider themselves hockey fans. You randomly select 9 of the population and ask each if they are hockey fans. Construct a binomial distribution and use the distribution to find the mean, variance, and standard deviation of the binomial distribution.
Answer by stanbon(48568)

Suppose that twenty-eight percent of women consider themselves hockey fans. You randomly select 9 of the population and ask each if they are hockey fans. Construct a binomial distribution and use the distribution to find the mean, variance, and standard deviation of the binomial distribution.
-----------
Hopefully you have a calculator and can figure out the distribution
using binom(n,p,x) for n=9, p=0.28, x = 1, then 2, then 3,...,then 9
-------------
mean = np = 9*0.28
variance = npq = 9*0.28*0.72
standard = sqrt(npq) = sqrt(variance)
============
Cheers,
Stan H.